# SynthFormer: Equivariant Pharmacophore-based Generation of Molecules for Ligand-Based Drug Design

### v1

@@ 3.2.2 Data generation
Equivariant pharmacophore encoder In our approach, we represent pharmacophores through a one-hot encoding scheme. Specifically, we denote a pharmacophore feature vector as $\mathbf{f}=[f_{1},f_{2},\ldots,f_{n}]$, where each $f_{i}\in\{0,1\}$ and $\sum_{i=1}^{n}f_{i}=1$. This representation allows us to capture the presence or absence of distinct pharmacophoric features in a binary format. Additionally, we employ Cartesian coordinate features to preserve equivariance. We treat this as a fully connected graph, because our hypothesis space is not limited by any bonds. We use the EGNN (Satorras et al., 2021), that takes as input the set of node embeddings $\mathbf{h}=\{\mathbf{h}_{0},\ldots,\mathbf{h}_{M-1}\}$, coordinate embeddings $\mathbf{x}^{\prime}=\{\mathbf{x}_{0}^{\prime},\ldots,\mathbf{x}_{M-1}^{\prime}\}$, and edge information $\mathcal{E}=(e_{ij})$ and outputs a transformation on $\mathbf{h}^{l+1}$ and $\mathbf{x}^{\prime l+1}$. The operation $\mathbf{h}^{l+1},\mathbf{x}^{\prime l+1}=\text{EGCL}(\mathbf{h},\mathbf{x}^{\prime},\mathcal{E})$ denotes an equivariant graph convolutional layer (EGCL), a specific variant of equivariant neural networks (EGNNs) as introduced in Satorras et al. (2021). The equations that define this layer are the following:
$m_{ij}$ $=\phi_{e}\left(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},\left\|\mathbf{x}_{i}^{\prime}-\mathbf{x}_{j}^{\prime}\right\|_{2}^{2},a_{ij}\right)$ (1)
@@
In equation 1, we input the relative squared distance between two coordinates $\left\|\mathbf{x}_{i}^{\prime}-\mathbf{x}_{j}^{\prime}\right\|_{2}^{2}$ into the edge operation $\phi_{e}$. Later, we use the representation $\mathbf{h}^{l+1}$ and pass it through cross-attention to the decoder. We employ a seven-layer EGNN as the encoder in our model.
@@
Synthetically accessible molecule decoder The input to the transformer begins with a token [START], followed by first BB, then P1. The input represents the last final product from the reactions. The model stops when the building block MLP predicts the end token [END]. An input token is a morgan fingerprint and start token , denoted by $[\texttt{fp},start]$, where $f_{p}\in\{0,1\}^{1024}$ is the Morgan fingerprint of length 1024 and radius 2 (Morgan, 1965). Building block tokens are converted into embeddings by an MLP. Moreover we use positional encodings to encode the position for each token.
$\mathbf{h}_{i}^{(0)}=\text{PE}(i)+\begin{cases}\bm{\epsilon}_{\text{start}},\text{type}(T_{i})=\text{START}\\ \text{MLP}_{\text{fp}}(p_{i}),\text{type}(T_{i})=\text{BB}\\ \end{cases}$ (5)
The decoder uses the standart multihead attention mechanism. However, at the output of the decoder, we take the multihead output pass and concatenate with the learnable molecular embeddings.
$\text{MultiHead}(Q,K,V)=\text{Concat}(\text{head}_{1},...,\text{head}_{h})W^{O}$ (6)
$\text{where head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}),$ (7)
where the projections are parameter matrices $W_{i}^{Q}\in\mathbb{R}^{d_{\text{model}}\times d_{k}},W_{i}^{K}\in\mathbb{R}^{d_{\text{model}}\times d_{k}},W_{i}^{V}\in\mathbb{R}^{d_{\text{model}}\times d_{v}},$ and $W^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$. The encoder is composed of a stack of $N=7$ identical layers and we employ $h=8$ attention heads. As we pass the input tokens through the last multi-head attention mechanism to obtain the representation $Z$. This representation serves as the basis for predicting the next building block $BB_{i+1}$, using a softmax applied to a linear transformation of $Z$, also refered as MultiHead(Q, K, V). The corresponding equation is given by:
$BB_{i+1}\sim\text{Softmax}(\text{Linear}(Z_{i}))$ (8)
@@
Next, we concatenate the representation $Z_{i}$ with the embeddings of the predicted building block $BB_{i+1}$ and apply a linear transformation followed by a softmax function to predict the reaction $R_{i+1}$. This process can be expressed as:
$R_{i+1}\sim\text{Softmax}(\text{Linear}(\text{Concat}(Z_{i},\text{Embedding}(BB_{i+1}))))$ (9)
@@
The model uses both the predicted building block $BB_{i+1}$ and the reaction $R_{i+1}$ to generate the product, which is then fed back into the model as input for the next iteration. This process continues iteratively, generating new building blocks and reactions, until the model predicts an end token for the building block, signaling the completion of the process.
@@ 3.3 Training
The building block loss function for building block fingerprint where $\ell$ is the length of the sequence :
$L_{\text{bb}}=\frac{1}{n}\sum_{j=0}^{l-1}\text{CE}(\hat{BB}_{i+1},BB_{i+1}),$ (10)
where CE is the cross entropy loss function, for reaction type:
$L_{\text{rxn}}=\frac{1}{m}\sum_{i=0}^{\ell-1}\text{CE}(\hat{r}_{i+1},r_{i+1}).$ (11)
The final loss function is the sum of the two terms: $L=L_{\text{bb}}+L_{\text{rxn}}$.
@@ 3.4 Inference
At inference time, we prepare the pharmacophore data and the start token as inputs. We generate tokens one by one. The transformer decoder first receives a start token and predicts building block BB1, followed by reaction 0. Next, it takes the fingerprint of BB1 as input and predicts BB2, followed by reaction 1, which is applied to generate product 1. Product 1 is then used to predict building block 3 and reaction 2, producing product 3. This process repeats until the end token is generated. Finally, an end token [END] terminates the process and the last molecule is returned as the product.
4 Results
PDBID Ref Dock Average Dock Gen Min Dock Gen Tanimoto Murcko Gobbi
1x8d -6.17 -6.32 ± 0.49 -8.39 0.06 0.06 0.20
1xbo -10.79 -7.22 ± 1.08 -10.07 0.12 0.12 0.39
2afw -3.85 -7.89 ± 0.91 -9.13 0.10 0.07 0.38
2aog -10.45 -8.75 ± 1.24 -11.62 0.07 0.00 0.30
2bt9 -6.68 -6.73 ± 0.62 -9.23 0.04 0.02 0.21
3coy -12.41 -7.14 ± 4.36 -11.26 0.10 0.08 0.43
3ga5 -9.68 -3.33 ± 1.61 -5.60 0.05 0.03 0.23
4q6d -7.41 -7.53 ± 1.54 -10.07 0.07 0.10 0.32
5fl4 -8.09 -7.52 ± 0.83 -9.28 0.10 0.10 0.37
5ka1 -7.67 -6.74 ± 1.07 -9.05 0.07 0.06 0.42
Table 1: The table shows reference docking energies (Ref Dock) and the mean and minimum docking energies (Dock Gen) for 100 generated molecules docked to the same protein. Tanimoto, Murcko, and Gobbi similarities indicate the averaged structural resemblance between the 100 generated molecules and the reference ligands.
4.1 Designing active compounds
We evaluate the model by selecting 10 compounds from different PDB entries and extract their pharmacophores. These pharmacophores are used as inputs to generate 100 new compounds per original ligand. We assess the structural similarities between the generated and original compounds, followed by redocking the generated compounds. The docking scores are then compared to those of the original ligands to evaluate how closely the generated compounds mimic the binding affinity and pose.
In this study, we evaluated the structural similarities between 100 generated molecules (See Appendix A Figure 4) and reference ligands for various PDB entries. The docking results reveal a range of binding affinities for the generated molecules, with mean docking energies varying from -3.33 to -8.75 kcal/mol (see Table 1 column 3). Notably, we can see that for the PDB entries 1x8d, 2bt9, 416d, 5fl4 the average docking energy and the reference docking energy are extremely close, while the smallest docking energy is always lower than the reference docking energy, except for 3ga5.
The calculated similarity metrics—Tanimoto, Murcko, and Gobbi—provide insights into the structural resemblance between the generated molecules and reference ligands. Tanimoto and Murko similarities ranged from 0.03 to 0.12 showing that the similarity is extremely low. On the other hand, Gobbi similarities ranged from 0.20 to 0.43, showing the 2d Pharmacophore similarity is greater, but this is far from being identical. Thus, these results highlight that this model is able to maintain 2D pharmacophore similarity better than any kind of fingerprint similarity. Moreover, the Gobbi similarity is a 2D similarity and we prompt the model with 3D representations, explaining why the similarity scores are not closer to 1.
@@ 4.1 Designing active compounds
Table 2: Properties of the generated molecules, featuring molecular weights of the generated molecules(MW) ranging from 305.96 g/mol to 339.02 g/mol, indicating a diverse array of molecular sizes. LogP values fall within the expected range according to the Ghose filter (typically -0.4 to 5.6), and we obtain reasonable QED scores that reflect the drug-likeness of the molecules.
We evaluate the generative model by selecting 10 molecules from PDB Bind and generate 100 compounds for each. The evaluation focuses on two key aspects. First, we perform a similarity analysis, comparing the generated molecules to the original ones by calculating similarity scores using Morgan, Murcko and Gobbi fingerprints. Second, we analyze the range of physicochemical properties, such as logp, druglikeness, synthetic accessibility, to assess the diversity and novelty of the generated compounds. This analysis helps gauge the relevance and diversity of the generated molecules, ensuring they maintain desirable properties while exploring the new chemical space.
4.2 Embedding exploration
Figure 3: The embedding of the selected query molecule is used as the reference, and the three closest molecules are identified using cosine similarity for qualitative assessment. As observed, these molecules preserve significant structural similarity to the reference.
@@ 4.2 Embedding exploration
To further investigate the structural similarity between building blocks, we leveraged the embeddings produced by our transformer model. Specifically, we extracted these embeddings and employed cosine similarity to quantify the closeness of different building blocks.
Cosine similarity is a measure of the angular distance between two vectors, providing insight into their relative orientation in the embedding space. By applying this metric to the embeddings, we identified pairs of building blocks that are closely aligned according to the model’s learned representations.
@@
Our analysis revealed that the building blocks exhibited a high degree of similarity, particularly in terms of preserving ring structures and the number of nitrogen atoms. The embeddings captured these structural features with remarkable fidelity, allowing us to observe that similar building blocks not only shared comparable ring configurations but also maintained consistent nitrogen counts.
Moreover, we took 100 building blocks sampled 10 most similar building blocks based on cosine similarity and computed the max similarity. 30% of building blocks had an average of 0.67 Tanimoto similarity. This result underscores the effectiveness of the transformer model in encoding structural nuances and highlights its capacity to differentiate and preserve essential chemical features within the embedding space. The preservation of rings and nitrogen counts suggests that the model has effectively learned to represent these critical components of molecular structures, leading to meaningful and accurate similarity measures.
4.3 Property Optimisation through genetic embedding modifications
We implement a genetic algorithm approach for molecular optimization within a synthetic reaction tree framework, aiming to discover molecules with improved properties. In this part we enhance molecular properties by altering reagents in the reaction steps within the known domain of the hypothesis. For each of the 10 molecules in the Protein Data Bank (PDB), we apply these modifications, using cosine similarity to adjust the reagents in the predicted list to obtain the desired molecules and we resample the reactions. This process helps identify pathways that yield the most optimal molecules based on targeted properties.
Once modifications are applied, we re-dock the resulting molecules, compare their properties, such as binding affinity and physicochemical properties (e.g., logP, druglikeness), and analyze the distributions. By iterating through cycles of modification and evaluation, we quantify improvements in properties over 3 cycles. We generate new molecules score them pick the top molecules and repeat this process 3 times. We show that we can decrease the logp by 0.21 and for a group of molecules druglikeness increase 0.09, while the average energy increases by 0.03 kcal/mol (worse).
4.4 Hit Expansion
@@
Our model is also effective for hit expansion in drug discovery, providing a strategic method for identifying structurally similar and synthesizable analogs of hit molecules (Keseru & Makara, 2006; Levin et al., 2023). To demonstrate this, we applied our model to design inhibitors for all 10 proteins that have been used for the active ligand design task, following the setup from Levin et al. (2023).We constrained the generation process by initializing with the original structure and then sampling further for possible modifications. Our objective was to identify potentially active synthesizable molecules that are close to the molecule from the crystal structure.

### v2

@@ 3.3 Architecture
Equivariant Pharmacophore Encoder We employ Cartesian coordinate features to preserve equivariance with the pharmacophore representation p to encode the pharmacophoric profile. We treat this as a fully connected graph, because our hypothesis space is not limited by any bonds. We use the EGNN (Satorras et al.,, 2021), that takes as input the set of node embeddings $\mathbf{p}=\{\mathbf{p}_{0},\ldots,\mathbf{p}_{M-1}\}$, coordinate embeddings $\mathbf{x}^{\prime}=\{\mathbf{x}_{0}^{\prime},\ldots,\mathbf{x}_{M-1}^{\prime}\}$, and edge information $\mathcal{E}=(e_{ij})$ and outputs a transformation on $\mathbf{h}^{l+1}$ and $\mathbf{x}^{\prime l+1}$. The equations that define this layer are the following:
$m_{ij}=\phi_{e}\left(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},\left\|\mathbf{x}_{i}^{\prime}-\mathbf{x}_{j}^{\prime}\right\|_{2}^{2},e_{ij}\right)$ (1)
@@
In Eq. 1 , we input the relative squared distance between two coordinates $\left\|\mathbf{x}_{i}^{\prime}-\mathbf{x}_{j}^{\prime}\right\|_{2}^{2}$ into the edge operation $\phi_{e}$. Later, we use the representation $\mathbf{h}^{l+1}$ and pass it through cross-attention to the decoder.Finally, $\varphi_{e}$ and $\varphi_{h}$ are the edge and node operations, respectively, which are commonly approximated by Multilayer Perceptrons (MLPs). We employ a seven-layer EGNN as the encoder in our model.
@@
Synthesis-Aware Decoder The input to the transformer begins with a token $[START]$, followed by the first B, then $P_{1}$. The input represents the last final product from the reactions. The model stops when the building block MLP predicts the end token $[END]$. An input token is a Morgan fingerprint and start token, denoted by $[f_{p},start]$, where $f_{p}\in\{0,1\}^{4096}$ is the Morgan fingerprint of length 4096 and radius 3 (Morgan,, 1965). Morgan fingerprint encodes local atomic environments by applying a circular algorithm that captures patterns in the arrangement and connectivity of atoms within a specified radius around each atom. This algorithm considers features such as atomic types, bonding patterns, and the functional groups in the neighborhood of each atom. The result is a fixed-length bit vector that summarizes these structural patterns. Building block tokens are converted into embeddings by an MLP. Moreover, we use positional encodings (PE) to encode the position for each molecule. The decoder uses the standard multihead attention mechanism. However, at the output of the decoder, we take the multihead output pass and concatenate with the learnable molecular embeddings.
$\text{MultiHead}(Q,K,V)=\text{Concat}(\text{h}_{1},...,\text{h}_{h})W^{O}$ (1)
$\text{where h}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}),$ (2)
and the projections are parameter matrices $W_{i}^{Q}\in\mathbb{R}^{d_{\text{model}}\times d_{k}},W_{i}^{K}\in\mathbb{R}^{d_{\text{model}}\times d_{k}},W_{i}^{V}\in\mathbb{R}^{d_{\text{model}}\times d_{v}},$ and $W^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$. The encoder is composed of a stack of $N=7$ identical layers and we employ $h=8$ attention heads. We pass the input tokens through the last multi-head attention mechanism to obtain the representation Z. During training, we utilize cosine similarity to guide the model towards generating molecules with desired properties. This involves calculating the cosine similarity between the generated molecule’s representation ($Z^{\prime}$) and the target molecule’s representation ($Z$), and using this similarity score as a component of the loss function. During sampling, we leverage the trained model to generate new molecules by maximizing the cosine similarity between the generated molecule’s representation and the representation embeddings, effectively guiding the generation process towards molecules that align with the learned chemical space:
$B_{i+1}=argmax(\text{Similarity}\left(Z_{i},Z^{\prime})\right))$ (3)
@@
Next, we concatenate the representation $Z_{i}$ with the embeddings of the predicted building block $B_{i+1}$ and apply a linear transformation followed by a softmax function to predict the reaction $R_{i+1}$. This process can be expressed as:
$R_{i+1}\sim\text{Softmax}(\text{Linear}(\text{Concat}(Z_{i},\text{f}_{p}(B_{i+1}))))$ (4)
@@
The model uses both the predicted building block $B_{i+1}$ and the reaction $R_{i+1}$ to generate the product, which is then fed back into the model as input for the next iteration. This process continues iteratively, generating new building blocks and reactions, until the model predicts an end token for the building block, signaling the completion of the process.
@@ 3.4 Training
The building block loss function for the building block fingerprint, where $\ell$ is the length of the sequence, is defined:
$L_{\text{B}}=\frac{1}{n}\sum_{j=0}^{l-1}\frac{Z^{\prime}_{i+1}\cdot Z_{i+1}}{||Z^{\prime}_{i+1}||\cdot||Z_{i+1}||}$ (5)
where $Z$ and $Z^{\prime}$ are vectors encoding the $Z$ representation and the encoded building block representation, respectively. The encoded Morgan fingerprint representation $Z^{\prime}$ is given:
$Z^{\prime}=\mathbf{W}\cdot\text{f}_{p}(B_{i+1})+\mathbf{b}$ (6)
The reaction loss function $L_{\text{rxn}}$ is defined:
$L_{\text{rxn}}=\frac{1}{m}\sum_{i=0}^{\ell-1}\text{CE}(\hat{r}_{i+1},r_{i+1})$ (7)
The final loss is the sum of the two terms defined:
$L=L_{\text{B}}+L_{\text{rxn}}$ (8)
@@ 3.5 Inference
At inference time, we prepare the pharmacophore data and the start token as inputs. The transformer decoder first receives a start token and predicts building block $B_{1}$, followed by reaction $R_{0}$. Next, it takes the fingerprint of $B_{1}$ as input and predicts $B_{2}$, followed by reaction $R_{1}$, which is applied to generate product $P_{1}$. Product $P_{1}$ is then used to predict building block $B_{3}$ and reaction $R_{2}$, producing product $P_{3}$. This process repeats until the end token is generated. Finally, an end token $[END]$ terminates the process and the last molecule is returned as the product.
4 Experiments
In this section, we mainly examine the performance of our model in the context of active ligand design. First, we compare the model against a random baseline to assess its ability to generate biologically relevant molecules. Next, we evaluate its performance against synthesis-aware models, highlighting its advantages in generating molecules that are diverse and active. We further compare our approach with three ligand-based 3D models, focusing on its docking performance and its unique ability to generate 100% synthesisable molecules. Moving beyond ligand design, we explore the utility of embedding encodings, by generating new analogs with different properties. Finally, we demonstrate the model’s potential for molecular optimization.
4.1 Experimental Setup
Reaction Templates We use the publicly available reaction template set (Hartenfeller et al.,, 2012). The template set contains 58 reaction templates encoded in SMARTS format.
Building Blocks We use commercially available building blocks (BBs) from Enamine (Ltd.,, 2025), which are molecules prepared in bulk to be readily synthesised into candidate molecules. There are 251,222 building blocks.
Similarity Scores To evaluate the similarity between the input molecule and the output molecule, we use the Tanimoto similarity score on three different fingerprints: (1) Morgan fingerprint of length 4096 and radius 2 (Morgan,, 1965), (2) Morgan fingerprint of Murcko scaffold (Bemis and Murcko,, 1996). The three similarity scores indicate chemical similarities in three different aspects: overall structure, scaffold structure, and pharmacophore property, respectively, and are all normalized to [0, 1].
Docking We used SMINA (Koes et al.,, 2013) for docking, converting ligands and proteins to PDBQT format with OpenBabel (O’Boyle et al.,, 2011). The docking box, centered on each ligand’s centroid, was set to 25x25x25 Å. Up to 10 poses were generated for each ligand docked into the protein’s binding site. This process was repeated for all ligand-protein pairs for further analysis.
4.2 Main Evaluation: Designing Active Compounds
Structural similarities between the generated and original compounds were assessed, followed by redocking of the generated compounds. Docking scores were compared to those of the original ligands to evaluate binding affinity and pose. Notably, for PDB entries 1x8d, 2bt9, 416d, and 5fl4, the average docking energy closely matched the reference, with the lowest docking energy consistently outperforming the reference, except for 3ga5 (see Table 2). Mean docking energies for generated molecules ranged from -4.55 to -10.14 kcal/mol (Appendix A to see similar molecules, Figure 3).
@@ 4.2 Main Evaluation: Designing Active Compounds
Table 1: The generated molecules exhibit diverse molecular weights (305.96–339.02 g/mol), LogP values within the Ghose filter range (-0.4 to 5.6), and reasonable drug-likeness (QED).
[…]
1x8d -6.17 -6.43 (-8.33) -4.62 0.07 0.10
1xbo -10.79 -8.22 (-11.22) -3.75 0.11 0.12
2afw -3.85 -7.92 (-9.21) -4.96 0.08 0.09
2aog -10.45 -8.63 (-11.53) -4.13 0.01 0.07
2bt9 -6.68 -7.63 (-10.11) -5.72 0.03 0.08
3coy -12.41 -10.14 (-12.26) -4.08 0.07 0.08
3ga5 -9.68 -4.55 (-7.60) -4.62 0.04 0.05
4q6d -7.41 -7.55 (-10.07) -3.52 0.10 0.07
5fl4 -8.09 -8.01 (-9.44) -3.94 0.12 0.11
5ka1 -7.67 -6.44 (-9.71) -5.21 0.07 0.07
Table 2: The table presents docking energies of the reference compound from the PDBID (Ref Dock) and the mean and minimum docking energies (Av. Dock Gen (Min)) for 100 generated molecules docked to the same protein as well as the Random Baseline from Enamine space. Tanimoto and Murcko similarities reflect the average resemblance between the 100 generated molecules and the reference ligands.
The similarity metrics, Tanimoto and Murcko, provide insights into the structural resemblance between the generated molecules and reference ligands indicating extremely low similarity, ranging from 0.06 to 0.12. As shown in Table 1, the generated molecules adhere to the Ghose filter, with molecular weights ranging from 305.96 to 339.02 g/mol, LogP values within the acceptable range (-0.4 to 5.6), and reasonable drug-likeness.
[…]
ChemProjector Synthesis 2.55 0.53 100%
SynNet Synthesis 2.7 0.64 100%
SQUID 3D 2.39 0.24 23.6%
Ligdream 3D 2.78 0.22 32.4%
SynthFormer Synthesis+3D 2.46 0.09 100%
Table 3: Comparison of synthesis-aware, 3D generative methods and SynthFormer for generating molecules across 10 PDB (100 compounds per PDB). SynthFormer achieves the lowest similarity to reference ligands while maintaining docking performance comparable to 3D generative models like SQUID. While, it generates more diverse and more active compounds than ChemProjector and SynNet.
Comparison to Random Baseline
In our experiments, as shown in Table 2, the best generated molecule with the best bose consistently outperforms the reference molecule in terms of docking energy. This result is particularly notable as we observe lower docking energies for the generated molecules across all cases compared to the reference, highlighting the superior performance of our approach. Furthermore, when compared to the random baseline, our generated molecules consistently beat the baseline in docking energy, indicating that our method is robust and effective in generating molecules with enhanced docking properties. The random baseline was achieved by sampling molecules from the available chemical space (Enamine building blocks and available reactions) and docking them to the binding sites. These results underline the efficacy of our model in optimizing molecular design for better binding performance.
Comparison Against Existing Methods
Table 3 presents a comparative analysis of molecule generation using synthesis models (ChemProjector (Luo et al.,, 2024) and SynNet (Gao et al.,, 2022)), ligand-based 3D models (LigDream (Skalic et al.,, 2019) and SQUID (Adams and Coley,, 2023)), and SynthFormer, for the task of generating 100 molecules for each target. ChemProjector is inherently designed for molecule-to-molecule tasks, which naturally results in the generation of analogs. The same applies to SynNet, making direct comparisons less fair. In contrast, 3D methods are guided only by scoring functions, and as demonstrated by SynFlowNet, the SA score is not a discriminatory feature for molecules. Nonetheless, this underscores the importance of SynthFormer as a model specifically tailored for synthesizable ligand-based design.
Synthesis scores for 3D models were taken from Drug Pose (Jocys et al.,, 2024), where they queried Enamine Real Space for existing molecules. For other models, the scores were based on the number of generated molecules with available synthetic pathways.
Luo et al. demonstrate strong performance in generating molecules similar to known actives, as evidenced by the low $\Delta$ Dock (2.55) and Tanimoto (0.53) scores.$\Delta$Dock (2.55) represents the average difference between the docking score of the reference compound and the docking scores of the generated compounds. These results suggest the generated molecules closely resemble existing drugs and exhibit favorable binding affinities. This observation highlights a limitation of Chemprojector and Synnet in a drug discovery context, where identifying molecules with new scaffolds is crucial for expanding the pool of potential drug candidates. SynthFormer, on the other hand is able to produce similarly active compounds as Chemprojector with low similarity.
In comparison to 3D methods, SynthFormer achieves a $\Delta$ Dock score of 2.46, which is slightly better than Ligdream (2.78) and comparable to SQUID (2.39), suggesting that SynthFormer generates molecules with docking performance similar to existing methods. In terms of chemical diversity, SynthFormer outperforms both SQUID and Ligdream with a significantly lower $\Delta$ Tanimoto score of 0.09, compared to SQUID (0.24) and Ligdream (0.22). This indicates that the molecules generated by SynthFormer are more distinct from known compounds. Furthermore, SynthFormer achieves a synthesisability of 100%, far exceeding the synthesisability of SQUID (23.6%) and Ligdream (32.4%), demonstrating that all molecules generated by SynthFormer are synthetically feasible. These results highlight the advantages of SynthFormer in generating diverse, chemically distinct, and synthetically accessible molecules.
4.3 Other Evaluations
We present additional capabilities of SynthFormer, that allow hit expansion and compound optimization, and we show evidence of meaningful encoding space organization due to effective training. None of these tasks have existing benchmarks for this particular framework and thus we do not compare against other compounds.
@@ 4.3 Other Evaluations
Building Block Encoding Exploration. To further investigate the structural similarity between building blocks, we leveraged the building block encodings produced in our synthesis aware decoder module denoted as Z’. Specifically, we extracted these encodings and used cosine similarity to quantify the closeness of different building blocks.
@@
We revealed that the encodings captured these structural features with remarkable fidelity, observing that similar building blocks not only shared comparable ring configurations but also maintained consistent atom type counts.
Moreover, we selected 100 building blocks and identified the 10 most similar ones based on cosine similarity (see Appendix A, Figure 2), finding that top 30% of the building blocks had an average Tanimoto score of 0.67. This shows that the synthesis-aware decoder effectively captures structural nuances while preserving essential chemical features.
@@
Hit Expansion. Our model is also effective for hit expansion in drug discovery, providing a strategic method for identifying structurally similar and synthesizable analogs of hit molecules (Keseru and Makara,, 2006; Levin et al.,, 2023). To demonstrate this, we applied our model to design inhibitors for all 10 proteins that have been used for the active ligand design task, following the setup from Levin et al., (2023). We constrained the generation process by initializing with the original structure and then sampling further for possible modifications. Our objective is to identify potentially active synthesizable molecules that are close to the molecule from the crystal structure.
@@
Molecule Optimization. We implement a genetic algorithm approach (See Appendix B) for molecular optimization within a synthetic reaction tree framework, aiming to discover molecules with improved properties. In this part we enhance molecular properties by altering reagents in the reaction steps. For each of the 10 molecules in the Protein Data Bank (PDB), we generate new molecules and using cosine similarity we adjust the reagents in the predicted list to obtain the desired molecules and we resample the reactions. This process helps identify pathways that yield the most optimal molecules based on targeted properties.
Once modifications are applied, we re-dock the resulting molecules, compare their properties, such as binding affinity and physicochemical properties (e.g., logP, druglikeness). By iterating through cycles of modification and evaluation, we quantify improvements in properties over 3 cycles. We generate new molecules, score them, pick the top molecules, and repeat this process 3 times. We show that we can decrease the logp by 0.21 and for a group of molecules druglikeness increase 0.09, while the average energy increases by 0.03 kcal/mol (worse).